### v1

# Gregorio Curello Mannheim Ludvig Sinander

[…]
The comparative statics of dominancethanks: We are grateful for comments from Ilia Krasikov, Giacomo Lanzani, Eddie Schlee, and a seminar audience at Arizona State University. Martín Grunwald provided excellent research assistance. Curello acknowledges support from the German Research Foundation (DFG) through CRC TR 224 (Project B02).
Gregorio Curello
Mannheim
   Ludvig Sinander
Oxford
   Mark Whitmeyer
Arizona State
(17 December 2025)
Zusammenfassung
In finite problems comprising objects, situations, and an object- and situation-contingent payoff function, we study the comparative statics of the set of undominated objects, meaning those for which there exists no mixture over objects that is superior whatever the situation. We consider both weak and strict dominance (corresponding to different degrees of ‘strictness’ in the definition of superiority). Our main theorem characterises those payoff transformations which robustly expand the not-weakly-dominated and not-strictly-dominated sets: the necessary and sufficient condition is that payoffs be transformed separately across situations, in either a monotone–concave or a constant manner. We apply our results to Pareto frontiers and games.
1 Introduction
Undominatedness is a fundamental solution concept in economics. Abstractly, in a finite problem comprising a finite set of objects, a finite set of situations, and an object- and situation-dependent payoff function, the not-weakly-dominated (NWD) and not-strictly-dominated (NSD) sets comprise those objects which are not weakly/strictly dominated by any mixture over objects, where ‘weakly dominated’ means weakly superior in every situation and strictly superior in some, and ‘strictly dominated’ means strictly superior in every situation. In a given player’s action-choice problem in a normal-form game, the objects are her own actions, the situations are profiles of actions of the other players, and weak and strict dominance coincide with the familiar game-theoretic concepts bearing those names. In a social planner’s problem, the objects are allocations, the situations are individuals, payoffs describe each individual’s preferences over (lotteries over) allocations, and the NWD set is known as the Pareto frontier.
In this paper, we study the comparative statics of dominance: which transformations of payoffs expand or contract the undominated sets? We are inspired by [Weinstein2016, BattigalliEtal2016], who identified sufficient conditions for expansion of the NSD set; we seek conditions that are necessary and sufficient.
Following the comparative-statics literature [[, e.g.]]Topkis1978,MilgromShannon1994,QuahStrulovici2009, we formalise our questions in a robust, detail-free way, by asking which payoff transformations expand the NWD set in every finite problem, and similarly for NSD and for contraction (rather than expansion) of the NWD and NSD sets.
Our main theorem provides the answer: a payoff transformation robustly contracts the NWD and NSD sets if and only if it is situation-wise, i.e. transforms payoffs separately situation by situation, and each situation’s transformation is convex and strictly increasing. Similarly, a payoff transformation robustly expands the NWD (NSD) set if and only if it is situation-wise and either (i) each situation’s transformation is concave and strictly (weakly) increasing or (ii) the transformation is degenerate in a precise sense to do with constant transformations. Economically, situation-wiseness means that within-situation payoffs are never contaminated by payoffs in other situations, and (by Pratt’s theorem,) convexity/concavity plus strict monotonicity means precisely decreased/increased risk-aversion in each situation.
Our above definition of dominance is the standard one whereby an object is undominated only if no mixture over objects dominates it. In an extension, we characterise the comparative statics of the sets of objects that are not (weakly or strictly) dominated by any object. The conditions are weaker: monotonicity remains necessary, but concavity/convexity does not, and situation-wiseness remains necessary for expansion but not for contraction.
Our first application is to Pareto dominance. Here, the objects $X$ are allocations, the situations $\{1,2,\dots,n\}$ are individuals, $U(\cdot,i):X\to\mathbf{R}$ is the payoff function of individual $i$, and the not-weakly-dominated set is called the Pareto frontier. Our main question in this application is which transformations of individuals’ payoffs expand or contract the Pareto frontier in every sub-economy (featuring only a subset of individuals). By our main theorem, the answer is (approximately): separate, individual-by-individual increases or decreases of risk-aversion.
Our second application is to strategic dominance in games. We fix a non-empty finite set $I$ of players, and consider all finite normal-form games $(I,(A_{i},u_{i})_{i\in I})$ played by $I$. For a given player $i\in I$, the objects $A_{i}$ are her own actions, the situations $A_{-i}$ are profiles of her opponents’ actions, her payoff function is $u_{i}:A_{i}\times A_{-i}\to\mathbf{R}$, and weak and strict dominance coincide with the usual game-theoretic concepts bearing those names. Our main theorem characterises those payoff transformations which robustly expand or contract the not-weakly-dominated and not-strictly-dominated sets. Applying this result repeatedly yields comparative statics for the set of rationalisable action profiles.
1.1 Related literature
The classic literature on undominatedness (also known as ‘admissibility’) spans statistical decision theory, game theory, and social choice. Much of the focus of this literature is on identifying conditions under which every undominated object is ‘sometimes-optimal’, i.e. expected-payoff-maximising against some (subjective) probability distribution over situations. [^1] See [Wald1939, Stein1955, fishburn1975separation, Pearce1984, Borgers1993, Battigalli1997, BattigalliEtal2016, CheKimKojimaRyan2024, ChengBorgers2025]. We ask a different sort of question, concerning comparative statics.
Two beautiful theorems in [CheKimKojima2021] identify those payoff transformations which lead the NWD set to ‘shift up’ (i.e. become ‘higher’) in a set-wise sense—specifically, in the weak set order. Our comparative-statics question is orthogonal to theirs: what we ask is which transformations of payoffs lead the NWD and NSD sets to expand (or contract).
Most closely related are [Weinstein2016, BattigalliEtal2016], who consider broadly the same comparative-statics question as we do, focusing on strict dominance. [^2] See also [[]section 8.1]deOliveiraLamba2025. What these authors show is that any situation-independent concave and strictly increasing transformation of payoffs expands the NSD set (we recover this result in section˜5.2 below). Our theorem extends and qualifies this finding, by identifying the entire class of payoff transformations which enjoy this expansion property, and also identifying those which enjoy the reverse ‘contraction’ property. We furthermore answer the same questions for the NWD set.
At a higher level, we contribute to the comparative-statics literature [[, e.g.]]Topkis1978,MilgromShannon1994,QuahStrulovici2009. The most closely related paper here is [QuahStrulovici2012], as we detail in a short companion paper [CurelloSinanderWhitmeyer2025note].
[…]
1.2 Roadmap
We rehearse the relevant background in section˜2. In section˜3, we introduce our setup. Section˜4 contains our main result, Theorem˜1, characterising which transformations robustly expand or contract the undominated sets. The applications to Pareto dominance and to games are studied in section˜5. Section˜6 contains a mirror theorem for dominance by objects rather than by mixtures (Theorem˜2). All proofs omitted from the main text may be found in the appendix.
2 Background
In this section, we tersely recount some of the theory of comparative risk-aversion, as well as a theorem of Sierpiński on concave functions.
2.1 Comparative risk aversion
Given a non-empty finite set $X$, we write $\Delta(X)$ for the set of all mixtures over $X$, i.e. all functions $p:X\to[0,1]$ which satisfy $\sum_{x\in X}p(x)=1$. Recall that an expected-utility decision-maker is one for whom there exists a function $u:X\to\mathbf{R}$ such that she weakly prefers a mixture $p\in\Delta(X)$ to another mixture $q\in\Delta(X)$ if and only if $\sum_{x\in X}u(x)p(x)\geq\sum_{x\in X}u(x)q(x)$.
Definition 1 ([Yaari1969]).
For a non-empty set $X$ and functions $u,v:X\to\mathbf{R}$, we say that $v$ is more risk-averse than $u$ if and only if for any $x\in X$ and $p\in\Delta(X)$, $\sum_{y\in X}v(y)p(y)\geq\mathrel{(>)}v(x)$ implies $\sum_{y\in X}u(y)p(y)\geq\mathrel{(>)}u(x)$.
[…]
Pratt’s theorem.
For a non-empty finite set $X$ and functions $u,v:X\to\mathbf{R}$, the following are equivalent:
  1. (a)
    $v$ is more risk-averse than $u$.
  2. (b)
    There exists a concave and strictly increasing function $\phi:\operatorname*{co}(u(X))\to\mathbf{R}$ such that $v(x)=\phi(u(x))$ for every $x\in X$.
  3. (c)
    The following two properties hold:
    1. (i)
      For any $x,y\in X$, $u(x)\geq\mathrel{(>)}u(y)$ implies $v(x)\geq\mathrel{(>)}v(y)$.
    2. (ii)
      For any $x,y,z\in X$, if $u(x)<u(y)<u(z)$, then
      $\frac{u(z)-u(y)}{u(y)-u(x)}\geq\frac{v(z)-v(y)}{v(y)-v(x)}.$
Remark 1.
In the literature, comparative risk-aversion and its characterisation are almost only ever considered in the special case in which the objects are monetary: $X\subseteq\mathbf{R}$. However, by inspection, properties (a)(c) are equally meaningful whatever the nature of the objects, and in fact they are equivalent even outside of the monetary case, as asserted above. See [CurelloSinanderWhitmeyer2025] for a proof.
When considering multiple situations $\mathcal{S}$ and two payoff functions $U,V:X\times\mathcal{S}\to\mathbf{R}$, Pratt’s theorem may be applied situation by situation: for each $s\in\mathcal{S}$, apply the theorem to $u(\cdot)\coloneqq U(\cdot,s)$ and $v(\cdot)\coloneqq V(\cdot,s)$.
2.2 Sierpiński’s theorem on concave functions
Several of our proofs rely on the following beautiful result due to [Sierpinski1920].
Sierpiński’s theorem.
A Lebesgue-measurable function $f:\mathbf{R}\to\mathbf{R}$ is concave if and only if for all $x,y\in\mathbf{R}$, $\frac{1}{2}f(x)+\frac{1}{2}f(y)\leq f(\frac{1}{2}x+\frac{1}{2}y)$.
3 Setup
In this section, we introduce our environment, formalise the notion of a payoff transformation, and pose our comparative-statics question(s).
3.1 Environment
There is a non-empty finite set $\mathcal{S}$ of situations. A decision-maker faces a set $X$ of objects, worries about some subset $S\subseteq\mathcal{S}$ of situations, and earns payoff $U(x,s)\in\mathbf{R}$ from object $x\in X$ in situation $s\in\mathcal{S}$. Formally, we consider all finite problems, meaning tuples $(X,S,U)$ where $X$ is a non-empty finite set, $S$ is a non-empty subset of $\mathcal{S}$, and $U$ is a map $X\times\mathcal{S}\to\mathbf{R}$. [^4] A technical detail is that ‘all finite problems’ is not actually a set (in ZFC), since ‘all non-empty finite sets $X$’ is not a set (in ZFC). The way around this annoyance is to fix throughout an (arbitrary) infinite set $\mathcal{X}$ of imaginable objects, and to restrict the definition of a finite problem $(X,S,U)$ by additionally requiring that $X\subseteq\mathcal{X}$. To save clutter, we leave all such $\subseteq\mathcal{X}$’ caveats implicit throughout the paper. We contemplate random choice, and assume expected utility, so that the payoff from mixture $p\in\Delta(X)$ in situation $s\in\mathcal{S}$ is $\sum_{x\in X}U(x,s)p(x)$. [^5] Here $\Delta(X)$ denotes the set of all functions $p:X\to[0,1]$ satisfying $\sum_{x\in X}p(x)=1$.
The set of not-strictly-dominated objects in a finite problem $(X,S,U)$ is
$\text{NSD}(X,S,U)\coloneqq\bigl\{x\in X:{}$ there is no $p\in\Delta(X)$ such that
$\text{$\textstyle\sum_{y\in X}U(y,s)p(y)>U(x,s)$ for every $s\in S$}\bigr\},$
and the set of not-weakly-dominated objects is
[…]
Note that for each object $x\in X$ in a finite problem $(X,S,U)$, we may view $U(x,\cdot):\mathcal{S}\to\mathbf{R}$ as a vector in $\mathbf{R}^{\mathcal{S}}$ (namely, the vector $\mathopen{}\mathclose{{\left(U(x,s)}}\right)_{s\in\mathcal{S}}$). As usual, for vectors $u,v\in\mathbf{R}^{n}$ (where $n\in\mathbf{N}$), ‘$u>v$’ means $u\geq v$ and $u\neq v$’, and $u\gg v$ means that $u_{i}>v_{i}$ for all $i\in\mathopen{}\mathclose{{\left\{1,\dots,n}}\right\}$. [^6] Hence $\text{NSD}(X,S,U)=\{x\in X:\text{$\nexists p\in\Delta(X)$ s.t. $\textstyle\sum_{y\in X}U(y,\cdot)p(y)\gg U(x,\cdot)$}\}$ and $\text{NWD}(X,S,U)=\{x\in X:\text{$\nexists p\in\Delta(X)$ s.t. $\textstyle\sum_{y\in X}U(y,\cdot)p(y)>U(x,\cdot)$}\}$
3.2 Payoff transformations
For a given finite problem $(X,S,U)$, our question is which transformations of the payoff $U$ cause the undominated sets $\text{NSD}(X,S,U)$ and $\text{NWD}(X,S,U)$ to expand, and which ones cause them to contract. By a transformation, we mean a map $\Psi:\mathbf{R}^{X\times\mathcal{S}}\to\mathbf{R}^{X\times\mathcal{S}}$, which associates each payoff function $U:X\times\mathcal{S}\to\mathbf{R}$ with a(nother) payoff function $\Psi(U)]:X\times\mathcal{S}\to\mathbf{R}$.
We restrict attention to transformations $\Psi$ that treat objects neutrally and separately. The former property requires that for any payoff functions $U,V:X\times\mathcal{S}\to\mathbf{R}$ and any permutation $\pi:X\to X$ such that $U(x,\cdot)=V(\pi(x),\cdot)$ for all $x\in X$, we have $\Psi(U)](x,\cdot)=[\Psi(V)](\pi(x),\cdot)$ for all $x\in X$. The latter property requires that for any $U,V:X\times\mathcal{S}\to\mathbf{R}$ and $x\in X$ such that $U(x,\cdot)=V(x,\cdot)$, we have $\Psi(U)](x,\cdot)=[\Psi(V)](x,\cdot)$.
Recall that for each $x\in X$, we may view $U(x,\cdot)$ as a vector in $\mathbf{R}^{\mathcal{S}}$. It is easy to see that if a transformation $\Psi:\mathbf{R}^{X\times\mathcal{S}}\to\mathbf{R}^{X\times\mathcal{S}}$ treats objects neutrally and separately, then there exists a map $\Phi:\mathbf{R}^{\mathcal{S}}\to\mathbf{R}^{\mathcal{S}}$ such that $\Psi(U)=\Phi\circ U$ for any $U:X\times\mathcal{S}\to\mathbf{R}$. We henceforth identify each transformation $\Psi$ with its associated map $\Phi:\mathbf{R}^{\mathcal{S}}\to\mathbf{R}^{\mathcal{S}}$.
Each family $(\phi_{s})_{s\in\mathcal{S}}$ of functions $\mathbf{R}\to\mathbf{R}$ generates a transformation $\Phi:\mathbf{R}^{\mathcal{S}}\to\mathbf{R}^{\mathcal{S}}$ via $(u_{s})_{s\in\mathcal{S}}\mapsto(\phi_{s}(u_{s}))_{s\in\mathcal{S}}$. Such transformations are ‘situation-wise’: for each situation $s\in\mathcal{S}$, the situation-$s$ output $\phi_{s}(u_{s})\in\mathbf{R}$ depends only on the situation-$s$ input $u_{s}\in\mathbf{R}$. Formally, given a family $(\phi_{s})_{s\in\mathcal{S}}$ of functions $\mathbf{R}\to\mathbf{R}$, we say that a transformation $\Phi:\mathbf{R}^{\mathcal{S}}\to\mathbf{R}^{\mathcal{S}}$ is $(\phi_{s})_{s\in\mathcal{S}}$-situation-wise if and only if $\Phi((u_{s})_{s\in\mathcal{S}})=(\phi_{s}(u_{s}))_{s\in\mathcal{S}}$ for every $(u_{s})_{s\in\mathcal{S}}\in\mathbf{R}^{\mathcal{S}}$.
3.3 Comparative-statics question(s)
Our question is which payoff transformations $\Phi:\mathbf{R}^{\mathcal{S}}\to\mathbf{R}^{\mathcal{S}}$ cause the undominated sets to expand or contract. We seek a ‘robust’ answer to this question, which does not depend on which particular finite problem is considered. Concretely, we are interested in the following four properties which a given transformation $\Phi:\mathbf{R}^{\mathcal{S}}\to\mathbf{R}^{\mathcal{S}}$ may or may not enjoy:
  1. (1)
    $\text{NSD}(X,S,U)\subseteq\text{NSD}(X,S,\Phi\circ U)$ for every finite problem $(X,S,U)$.
  2. (2)
    $\text{NWD}(X,S,U)\subseteq\text{NWD}(X,S,\Phi\circ U)$ for every finite problem $(X,S,U)$.
  3. (3)
    $\text{NSD}(X,S,\Phi\circ U)\subseteq\text{NSD}(X,S,U)$ for every finite problem $(X,S,U)$.
  4. (4)
    $\text{NWD}(X,S,\Phi\circ U)\subseteq\text{NWD}(X,S,U)$ for every finite problem $(X,S,U)$.
For each of these four properties, our question is which transformations $\Phi$ enjoy that property.
By asking ‘robust’ questions, requiring comparative statics to hold across all finite problems, we obtain answers which are independent of the details of any given problem. The answers to our ‘robust’ questions will furthermore turn out to be simple, providing straightforward sufficient conditions for comparative statics which are as weak as possible (subject to ‘robustness’).
In more detail, the advantage of considering various sets $X$ of objects is that this yields conclusions which are robust to any constraints which the decision-maker may face. Similarly, considering different sets $S\subseteq\mathcal{S}$ of situations yields conclusions that are robust to the decision-maker having a particular model or world-view, in which only certain situations are contemplated. For example, the arrival of new information may allow the decision-maker to rule out certain situations.
In some applications, ruling out situations is not natural or well-motivated. For such cases, we will have some results below that are ‘partially robust’, considering only finite problems $(X,S,U)$ such that $S=\mathcal{S}$.
4 Main theorem
Theorem 1.
Assume that $2\leq\mathopen{}\mathclose{{\left\lvert\mathcal{S}}}\right\rvert$$\infty$, and consider a transformation $\Phi:\mathbf{R}^{\mathcal{S}}\to\mathbf{R}^{\mathcal{S}}$.
  1. (a)
    Property (1) holds if and only if there exists a family $(\phi_{s})_{s\in\mathcal{S}}$ of weakly increasing functions $\mathbf{R}\to\mathbf{R}$ such that $\Phi$ is $(\phi_{s})_{s\in\mathcal{S}}$-situation-wise and either (i) $\phi_{s}$ is constant for all but at most one $s\in\mathcal{S}$ or (ii) $\phi_{s}$ is concave for all $s\in\mathcal{S}$.
  2. (b)
    Property (2) holds if and only if there exists a family $(\phi_{s})_{s\in\mathcal{S}}$ of functions $\mathbf{R}\to\mathbf{R}$ such that $\Phi$ is $(\phi_{s})_{s\in\mathcal{S}}$-situation-wise and either (i) $\phi_{s}$ is concave and strictly increasing for each $s\in\mathcal{S}$ or (ii) $\phi_{s}$ is constant for every $s\in\mathcal{S}$.
  3. (c)
    Property (3) holds if and only if Property (4) holds, which holds if and only if there exists a family $(\phi_{s})_{s\in\mathcal{S}}$ of convex and strictly increasing functions $\mathbf{R}\to\mathbf{R}$ such that $\Phi$ is $(\phi_{s})_{s\in\mathcal{S}}$-situation-wise.
The proofs of the three parts (a)(c) are similar to each other. Sufficiency follows from a short elegant argument substantially due to [Weinstein2016, BattigalliEtal2016] (although these authors use it to establish a weaker claim). Focusing on part (c), this argument goes as follows. The condition in (c) implies, by Pratt’s theorem, that applying $\Phi$ makes the decision-maker less risk-averse in every situation, meaning (by definition) that whenever a mixture $p\in\Delta(X)$ is (strictly) preferred to an object $x\in X$ in a situation $s\in S$ of a finite problem $(X,S,U)$, the same is true in the transformed problem $(X,S,\Phi\circ U)$. Thus any object $x\in X$ which was dominated (by some mixture $p\in\Delta(X)$) in $(X,S,U)$ remains dominated (by that same mixture $p$) in $(X,S,\Phi\circ U)$; in other words, applying $\Phi$ expands the dominated set, so shrinks the undominated set.
Necessity is more difficult. For each of the three parts (a)(c), we proceed in two steps, first showing that the transformation $\Phi$ must be situation-wise, and then that the functions $(\phi_{s})_{s\in\mathcal{S}}$ must satisfy the stated properties.
In the first step, there is a perhaps surprising asymmetry between expansion (parts (a) and (b)) and contraction (part (c)). It is almost immediate that the expansive properties (1) and (2) necessitate situation-wiseness. [^7] For any situation $s\in\mathcal{S}$, whenever $u,v\in\mathbf{R}^{\mathcal{S}}$ satisfy $u_{s}=v_{s}$, it must be that $\Phi_{s}(u)=\Phi_{s}(v)$, since otherwise letting $x\neq y$, $U(x,\cdot)\coloneqq u$ and $U(y,\cdot)\coloneqq v$, we have $\text{NSD}(\{x,y\},\{s\},U)=\{x,y\}\nsubseteq\text{NSD}(\{x,y\},\{s\},\Phi\circ U)$, and similarly for NWD.
By contrast, it takes some work to show that the contractive properties (3) and (4) necessitate situation-wiseness. It is easily shown that these properties necessitate strict situation-wise monotonicity: for any $u,v\in\mathbf{R}^{\mathcal{S}}$ and any situation $s\in\mathcal{S}$, $u_{s}>v_{s}$ implies $\Phi_{s}(u)>\Phi_{s}(v)$. [^8] If $\Phi_{s}(u)\leq\Phi_{s}(v)$ and $u_{s}>v_{s}$, then fixing $x\neq y$ and letting $U(x,\cdot)\coloneqq u$ and $U(y,\cdot)\coloneqq v$ yields $\text{NSD}(\{x,y\},\{s\},U)=\{x\}\nsupseteq\{y\}\subseteq\text{NSD}(\{x,y\},\{s\},\Phi\circ U)$, and similarly NWD. Having shown this, we next argue that for each $s\in\mathcal{S}$, the map $u_{s}\mapsto\Phi_{s}\mathopen{}\mathclose{{\left((u_{t})_{t\in\mathcal{S}}}}\right)$ must be continuous, by showing that any discontinuity would permit the construction of a finite problem $(X,\{s,t\},U)$ that violates properties (3) and (4). Together, strict situation-wise monotonicity and continuity imply situation-wiseness.
Having shown that properties (1)(4) each imply the existence of a family $(\phi_{s})_{s\in\mathcal{S}}$ of functions $\mathbf{R}\to\mathbf{R}$ such that the transformation $\Phi$ is $(\phi_{s})_{s\in\mathcal{S}}$-situation-wise, it remains to establish that $(\phi_{s})_{s\in\mathcal{S}}$ must have the properties asserted by parts (a)(c). This is accomplished by a pair of propositions given in the next section.
4.1 Situation-wise transformations
We now identify the conditions on situation-wise transformations that are necessary and sufficient for comparative statics. These results form part of the proof of Theorem˜1, but are also of independent interest.
We characterise not only the fully robust ‘for all finite problems’ comparative-statics properties (1)(4), but also the following four ‘partially robust’ comparative-statics properties in which the full set $\mathcal{S}$ of situations (rather than a subset $S\subseteq\mathcal{S}$) is always contemplated:
  1. (1enumi)
    $\text{NSD}(X,\mathcal{S},U)\subseteq\text{NSD}(X,\mathcal{S},\Phi\circ U)$ for every finite problem $(X,\mathcal{S},U)$.
  2. (2enumi)
    $\text{NWD}(X,\mathcal{S},U)\subseteq\text{NWD}(X,\mathcal{S},\Phi\circ U)$ for every finite problem $(X,\mathcal{S},U)$.
  3. (3enumi)
    $\text{NSD}(X,\mathcal{S},\Phi\circ U)\subseteq\text{NSD}(X,\mathcal{S},U)$ for every finite problem $(X,\mathcal{S},U)$.
  4. (4enumi)
    $\text{NWD}(X,\mathcal{S},\Phi\circ U)\subseteq\text{NWD}(X,\mathcal{S},U)$ for every finite problem $(X,\mathcal{S},U)$.
Characterising properties (1enumi)(4enumi) is valuable for applications, as we shall see in section˜5 below.
Our first proposition concerns only expansions of undominated sets.
Proposition 1.
Assume that $2\leq\mathopen{}\mathclose{{\left\lvert\mathcal{S}}}\right\rvert$$\infty$. Fix a family $(\phi_{s})_{s\in\mathcal{S}}$ of functions $\mathbf{R}\to\mathbf{R}$, and let $\Phi:\mathbf{R}^{\mathcal{S}}\to\mathbf{R}^{\mathcal{S}}$ be $(\phi_{s})_{s\in\mathcal{S}}$-situation-wise.
  1. (a)
    Property (1) holds if and only if either (i) $\phi_{s}$ is concave and weakly increasing for each $s\in\mathcal{S}$ or (ii) $\phi_{s}$ is weakly increasing for some $s\in\mathcal{S}$ and constant for all other $s\in\mathcal{S}$.
  2. (b)
    Property (1enumi) holds if and only if either (i) $\phi_{s}$ is concave and weakly increasing for each $s\in\mathcal{S}$ or (ii) $\phi_{s}$ is constant for some $s\in\mathcal{S}$.
  3. (c)
    Property (2) holds if and only if property (2enumi) holds, which holds if and only if either (i) $\phi_{s}$ is concave and strictly increasing for each $s\in\mathcal{S}$ or (ii) $\phi_{s}$ is constant for every $s\in\mathcal{S}$.
The sufficiency (‘if’) claims in Proposition˜1 follow from the ‘short elegant argument’ described immediately after Theorem˜1 above. The necessity of concavity and convexity follows from a kind of reversal of that argument; for part (c), it goes as follows. A failure of convexity means, by Pratt’s theorem, that applying the transformation $\Phi$ fails, in some situation $s\in\mathcal{S}$, to make the decision-maker less risk-averse. By definition, that means that there is a finite problem in which some object $x\in X$ is dis-preferred in situation $s$ to a mixture $p\in\Delta(X)$, but becomes preferred once the transformation is applied. Then (using Sierpiński’s theorem,) a finite problem can be constructed in which $x$ is dominated, but becomes undominated once the transformation is applied—so the undominated set fails to shrink.
Strict monotonicity is necessary in the NWD parts of Proposition˜1 because weak dominance requires ‘weakly better in every situation and strictly better in some’. Hence an object which is not weakly dominated may become weakly dominated when a weakly-but-not-strictly increasing transformation is applied. By contrast, concavity plus weak monotonicity are enough to rule out not-strictly-dominated objects from becoming strictly dominated (because strict dominance demands strict superiority in every situation). Similar logic explains the role of the constant maps.
Our second proposition concerns only contractions of undominated sets.
Proposition 2.
Assume that $2\leq\mathopen{}\mathclose{{\left\lvert\mathcal{S}}}\right\rvert$$\infty$. Fix a family $(\phi_{s})_{s\in\mathcal{S}}$ of functions $\mathbf{R}\to\mathbf{R}$, and let $\Phi:\mathbf{R}^{\mathcal{S}}\to\mathbf{R}^{\mathcal{S}}$ be $(\phi_{s})_{s\in\mathcal{S}}$-situation-wise. Property (3) holds if and only if property (3enumi) holds, which holds if and only if property (4) holds, which holds if and only if property (4enumi) holds, which holds if and only if $\phi_{s}$ is convex and strictly increasing for each $s\in\mathcal{S}$.
Proposition˜2 is similar to Proposition˜1, and is true for similar reasons. Strict monotonicity is required in Proposition˜2 since any weakly-but-not-strictly increasing transformation sometimes makes a previously dominated object undominated.
Remark 2.
By Pratt’s theorem, the condition in Proposition˜2 is equivalent to reducing risk-aversion in every situation. By another theorem of ours [[, Theorem 1 in]]CurelloSinanderWhitmeyer2025, reduced risk-aversion is behaviourally equivalent to the decision-maker gaining access to a (possibly random) outside option. It follows that dominated choices (in a decision problem or a game) can never be explained by the presence of an unobserved outside option: any choice that is dominated in the decision-maker’s problem exclusive of the outside option, $(X,S,U)$, is also dominated in the ‘true’ problem $(X,S,\Phi\circ U)$ that accounts for the outside option.
5 Applications
In this section, we apply our results to Pareto frontiers and to undominated (and rationalisable) actions in games.
5.1 Pareto dominance
Fix a population $\mathcal{S}\coloneqq\{1,2,\dots,n\}$ of $n\geq 2$ individuals. An economy is a pair $(X,U)$, where $X$ is a non-empty finite set of allocations and $U:X\times\mathcal{S}\to\mathbf{R}$ encodes payoffs: $U(\cdot,i):X\to\mathbf{R}$ is the payoff function of individual $i\in\mathcal{S}$. We assume expected utility, so that in an economy $(X,U)$, the payoff of individual $i\in\mathcal{S}$ from a mixture $p\in\Delta(X)$ is $\sum_{x\in X}U(x,i)p(x)$.
The Pareto frontier $\text{PF}_{\mathcal{S}}(X,U)$ of an economy $(X,U)$ is the set of all allocations such that no mixture is weakly preferred by all individuals and strictly preferred by some; in other words, $\text{PF}_{\mathcal{S}}(X,U)\coloneqq\text{NWD}(X,\mathcal{S},U)$. More generally, when only a (non-empty) sub-population $S\subseteq\mathcal{S}$ is considered, the Pareto frontier of economy $(X,U)$ is $\text{PF}_{S}(X,U)\coloneqq\text{NWD}(X,S,U)$.
Remark 3.
For any economy $(X,U)$ and sub-population $S\subseteq\mathcal{S}$, by the ‘cautious’ version of Pearce’s lemma [[, see e.g.]Theorem 1.7]Myerson1991, an allocation $x\in X$ belongs to $\text{PF}_{S}(X,U)$ if and only if there exist strictly positive weights $(\lambda_{i})_{i\in S}$ such that $x\in\operatorname*{arg\,max}_{y\in X}\sum_{i\in S}\lambda_{i}U(y,i)$. [^9] Curiously, this characterisation of NWD does not extend straightforwardly to infinite sets $X$ of objects/allocations [[, see]]CheKimKojimaRyan2024. By contrast, the familiar (‘incautious’) Pearce’s lemma for NSD [[, see e.g.]Theorem 1.6]Myerson1991 extends directly to infinite $X$ given some compactness and upper semi-continuity. In other words, membership of the Pareto frontier is equivalent to being the optimal choice of a utilitarian planner, for some Pareto weights. This motivates interest in the Pareto frontier as defined above (the deterministic allocations that are not dominated by any mixture over allocations). One could of course also consider an alternative definition in which only dominance by allocations is contemplated; we do this in section˜6 below.
Theorem˜1 characterises which payoff transformations generate expansion or contraction of the Pareto frontier $\text{PF}_{S}(X,U)$ of every economy $(X,U)$ for every sub-population $S\subseteq\mathcal{S}$. One upshot is that such transformations must be individual-by-individual, with no scope for one individual’s preferences to influence how another individual’s preferences are transformed. The other upshot (by Pratt’s theorem) is that contraction requires each individual to become less risk-averse, while (neglecting some exceptional cases) expansion requires that each individual become more risk-averse.
Depending on the application, the population may simply be fixed, in which case it is not natural to consider sub-populations. Propositions˜1 and 2 characterise, among individual-by-individual transformations, which ones expand or contract the full-population Pareto frontier $\text{PF}_{\mathcal{S}}(X,U)$ of every economy $(X,U)$. Again, the gist of the conditions is that expansion (contraction) requires individuals to become more (less) risk-averse.
5.2 Strategic dominance in games
Fix a non-empty set $I$ of players, with $\mathopen{}\mathclose{{\left\lvert I}}\right\rvert\geq 2$. We consider all finite normal-form games played by $I$, meaning all tuples $(I,(A_{i},u_{i})_{i\in I})$ such that for each $i\in I$, $A_{i}$ is a non-empty finite set and $u_{i}$ is a map $A\to\mathbf{R}$, where $A\coloneqq\prod_{j\in I}A_{j}$. The interpretation is that players simultaneously choose actions, and that for each player $i\in I$, her available actions are $A_{i}$ and her payoff is $u_{i}(a)$, where $a\in A$ is the profile of all players’ actions. Weak and strict dominance coincide with the familiar game-theoretic concepts bearing those names: player $i$’s not-weakly-dominated actions are $\text{NWD}(A_{i},A_{-i},u_{i})$, and her not-strictly-dominated actions are $\text{NSD}(A_{i},A_{-i},u_{i})$, where $A_{-i}\coloneqq\prod_{j\in I\setminus\{i\}}A_{j}$.
We contemplate general transformations $\Phi_{i}:\mathbf{R}^{A}\to\mathbf{R}^{A}$ of a player $i$’s payoff. By Theorem˜1, the transformations $\Phi_{i}$ which contract or expand $\text{NWD}(A_{i},A_{-i},u_{i})$ or $\text{NSD}(A_{i},A_{-i},u_{i})$ in every game $(I,(A_{i},u_{i})_{i\in I})$ must transform player $i$’s payoffs profile-by-profile: for any profile $a_{-i}\in A_{-i}$ of opponents’ actions, player $i$’s transformed payoffs $(\Phi_{i}\circ u_{i})(\cdot,a_{-i}):A_{i}\to\mathbf{R}$ can depend only on the pre-transformation payoffs $u_{i}(\cdot,a_{-i})$ against that same profile $a_{-i}\in A_{-i}$, not on the pre-transformation payoffs $u_{i}(\cdot,b_{-i})$ against any other profile $b_{-i}\in A_{-i}\setminus\{a_{-i}\}$. In terms of the payoff matrix of a two-player game, for the row player, this means a column-by-column transformation, with the transformed entries in each column depending only on the pre-transformed entries in that same column.
Theorem˜1 furthermore asserts that if a transformation is to contract player $i$’s undominated sets, then (by Pratt’s theorem) it must be that against each profile $a_{-i}\in A_{-i}$ of opponents’ actions, player $i$’s preferences over (mixtures over) her own actions becomes less risk-averse. Similarly, expansion requires—degenerate transformations aside—that player $i$ become more risk-averse, profile by profile.
For each player $i\in I$ and integer $n\in\mathbf{N}$, let $A_{i}^{n}\coloneqq\text{NSD}(A_{i}^{n-1},A_{-i}^{n-1},u_{i})$, where for each $i\in I$, $A_{-i}^{n}\coloneqq\prod_{j\in I\setminus\{i\}}A_{j}^{n}$ for all $n\in\mathbf{N}$, $A_{i}^{0}\coloneqq A_{i}$ and $A_{-i}^{0}\coloneqq A_{-i}$. The rationalisable action profiles are exactly $\prod_{i\in I}\bigcap_{n\in\mathbf{N}}A_{i}^{n}$. By applying Theorem˜1 iteratively for each $n\in\mathbf{N}$, we see that when all players become more (less) risk-averse in a profile-by-profile fashion, the set of rationalisable action profiles expands (shrinks).
A special case is when player $i$’s payoff is transformed in a profile-independent way: that is, $\Phi_{i}$ is not only $(\phi_{i}^{a_{-i}})_{a_{-i}\in A_{-i}}$-situation-wise, but there is a single function $\phi_{i}:\mathbf{R}\to\mathbf{R}$ such that $\phi_{i}^{a_{-i}}=\phi_{i}$ for every profile $a_{-i}\in A_{-i}$. In this special case, our comparative-statics questions become:
  1. (1)
    $\text{NSD}(A_{i},A_{-i},u_{i})\subseteq\text{NSD}(A_{i},A_{-i},\phi_{i}\circ u_{i})$ for every player $i\in I$ and every finite game $(I,(A_{j},u_{j})_{j\in I})$.
  2. (2)
    $\text{NWD}(A_{i},A_{-i},u_{i})\subseteq\text{NWD}(A_{i},A_{-i},\phi_{i}\circ u_{i})$ for every player $i\in I$ and every finite game $(I,(A_{j},u_{j})_{j\in I})$.
  3. (3)
    $\text{NSD}(A_{i},A_{-i},\phi_{i}\circ u_{i})\subseteq\text{NSD}(A_{i},A_{-i},u_{i})$ for every player $i\in I$ and every finite game $(I,(A_{j},u_{j})_{j\in I})$.
  4. (4)
    $\text{NWD}(A_{i},A_{-i},\phi_{i}\circ u_{i})\subseteq\text{NWD}(A_{i},A_{-i},u_{i})$ for every player $i\in I$ and every finite game $(I,(A_{j},u_{j})_{j\in I})$.
The following corollary of Propositions˜1 and 2 gives necessary and sufficient conditions for comparative statics in this special case. The ‘if’ half of part (1) recovers the result of [Weinstein2016, BattigalliEtal2016]; the rest is (presumably) new.
Corollary 1.
Assume that $2\leq\mathopen{}\mathclose{{\left\lvert I}}\right\rvert$$\infty$, and let $(\phi_{i})_{i\in I}$ be maps $\mathbf{R}\to\mathbf{R}$.
  1. (a)
    Property (1) holds if and only if for each player $i\in I$, $\phi_{i}$ is concave and weakly increasing.
  2. (b)
    Property (2) holds if and only if for each player $i\in I$, $\phi_{i}$ is either (i) concave and strictly increasing or (ii) constant.
  3. (c)
    Property (3) holds if and only if property (4) holds, which holds if and only if for each player $i\in I$, $\phi_{i}$ is convex and strictly increasing.
6 Dominance by objects (rather than mixtures)
Our results in the previous sections concerned the standard definition of ‘undominated’, which deems an object undominated only if it is not dominated by any mixture over objects. An alternative definition [[, in the spirit of]]Borgers1993 worries only about dominance by other objects:
$\widehat{\text{NSD}}(X,S,U)$ $\coloneqq\bigl\{x\in X:\text{there is no $y\in X$ such that}$
$\qquad\qquad\qquad\text{$U(y,s)>U(x,s)$ for every $s\in S$}\bigr\}$
$\text{and}\quad\widehat{\text{NWD}}(X,S,U)$ $\coloneqq\bigl\{x\in X:\text{there is no $y\in X$ such that}$
$U(y,s)\geq U(x,s)$ for every $s\in S$ and
$\qquad\qquad\qquad\text{$U(y,s)>U(x,s)$ for some $s\in S$}\bigr\}.$
In this section, we study comparative-statics for these alternative notions of undominatedness. Concretely, we consider the following properties which a given payoff transformation $\Phi:\mathbf{R}^{\mathcal{S}}\to\mathbf{R}^{\mathcal{S}}$ may or may not satisfy:
  1. ($\hskip 1.00006pt\widehat{1}\hskip 1.00006pt$)
    $\widehat{\text{NSD}}(X,S,U)\subseteq\widehat{\text{NSD}}(X,S,\Phi\circ U)$ for every finite problem $(X,S,U)$.
  2. ($\hskip 1.00006pt\widehat{2}\hskip 1.00006pt$)
    $\widehat{\text{NWD}}(X,S,U)\subseteq\widehat{\text{NWD}}(X,S,\Phi\circ U)$ for every finite problem $(X,S,U)$.
  3. ($\hskip 1.00006pt\widehat{3}\hskip 1.00006pt$)
    $\widehat{\text{NSD}}(X,S,\Phi\circ U)\subseteq\widehat{\text{NSD}}(X,S,U)$ for every finite problem $(X,S,U)$.
  4. ($\hskip 1.00006pt\widehat{4}\hskip 1.00006pt$)
    $\widehat{\text{NWD}}(X,S,\Phi\circ U)\subseteq\widehat{\text{NWD}}(X,S,U)$ for every finite problem $(X,S,U)$.
For each of these four properties, our question is which transformations $\Phi$ enjoy that property. The following theorem provides the answer.
Theorem 2.
Assume that $2\leq\mathopen{}\mathclose{{\left\lvert\mathcal{S}}}\right\rvert$$\infty$, and consider a transformation $\Phi:\mathbf{R}^{\mathcal{S}}\to\mathbf{R}^{\mathcal{S}}$.
  1. (a)
    Property ($\hskip 1.00006pt\widehat{1}\hskip 1.00006pt$) holds if and only if there exists a family $(\phi_{s})_{s\in\mathcal{S}}$ of weakly increasing functions $\mathbf{R}\to\mathbf{R}$ such that $\Phi$ is $(\phi_{s})_{s\in\mathcal{S}}$-situation-wise.
  2. (b)
    Property ($\hskip 1.00006pt\widehat{2}\hskip 1.00006pt$) holds if and only if there exists a family $(\phi_{s})_{s\in\mathcal{S}}$ of functions $\mathbf{R}\to\mathbf{R}$ such that $\Phi$ is $(\phi_{s})_{s\in\mathcal{S}}$-situation-wise and either (i) $\phi_{s}$ is strictly increasing for each $s\in\mathcal{S}$ or (ii) $\phi_{s}$ is constant for each $s\in\mathcal{S}$.
  3. (c)
    Property ($\hskip 1.00006pt\widehat{3}\hskip 1.00006pt$) holds if and only if for all $u,v\in\mathbf{R}^{\mathcal{S}}$ and $s\in\mathcal{S}$, $u_{s}>v_{s}$ implies $\Phi_{s}(u)>\Phi_{s}(v)$.
  4. (d)
    Property ($\hskip 1.00006pt\widehat{4}\hskip 1.00006pt$) holds if and only if for all $u,v\in\mathbf{R}^{\mathcal{S}}$ and $s\in\mathcal{S}$, both of the following hold: (i) $u_{s}>v_{s}$ implies $\Phi_{s}(u)>\Phi_{s}(v)$, and (ii) $u_{s}=v_{s}$ and $\Phi_{s}(u)>\Phi_{s}(v)$ together imply $u>v$.
The first two parts are straightforward to show, and the latter two are immediate. Note that situation-wiseness remains necessary for expansion, but not for contraction, as illustrated by the following example.
Example 1.
Let $\mathcal{S}=\{1,2\}$, and let $\Phi:\mathbf{R}^{2}\to\mathbf{R}^{2}$ be given by
[…]
By Theorem˜2, $\Phi$ satisfies properties ($\hskip 1.00006pt\widehat{3}\hskip 1.00006pt$) and ($\hskip 1.00006pt\widehat{4}\hskip 1.00006pt$). It is not situation-wise.
In spite of Theorem˜2, situation-wiseness is necessary for contraction when attention is restricted to moderately well-behaved transformations $\Phi$:
Observation 1.
Assume that $2\leq\mathopen{}\mathclose{{\left\lvert\mathcal{S}}}\right\rvert$$\infty$, and consider a continuous transformation $\Phi:\mathbf{R}^{\mathcal{S}}\to\mathbf{R}^{\mathcal{S}}$. Assume that for all $u,v\in\mathbf{R}^{\mathcal{S}}$ and $s\in\mathcal{S}$, $u_{s}>v_{s}$ implies $\Phi_{s}(u)>\Phi_{s}(v)$. Then there exists a family $(\phi_{s})_{s\in\mathcal{S}}$ of functions $\mathbf{R}\to\mathbf{R}$ such that $\Phi$ is $(\phi_{s})_{s\in\mathcal{S}}$-situation-wise and $\phi_{s}$ is strictly increasing for each $s\in\mathcal{S}$.
Beweis.
Fix any $u,v\in\mathbf{R}^{\mathcal{S}}$ and $s\in\mathcal{S}$ such that $u_{s}=v_{s}$; we must show that $\Phi_{s}(u)=\Phi_{s}(v)$. To that end, for each $\varepsilon>0$, define $\underline{u}^{\varepsilon},\overline{u}^{\varepsilon}\in\mathbf{R}^{\mathcal{S}}$ by $\underline{u}^{\varepsilon}_{s}\coloneqq u_{s}-\varepsilon$, $\overline{u}^{\varepsilon}_{s}\coloneqq u_{s}+\varepsilon$ and $\underline{u}^{\varepsilon}_{t}\coloneqq u_{t}\eqqcolon\overline{u}^{\varepsilon}_{t}$ for every $t\in\mathcal{S}\setminus\{s\}$. By the monotonicity assumption, $\Phi_{s}\mathopen{}\mathclose{{\left(\underline{u}^{\varepsilon}}}\right)<\Phi_{s}(v)<\Phi_{s}\mathopen{}\mathclose{{\left(\overline{u}^{\varepsilon}}}\right)$ for every $\varepsilon>0$. Since $\Phi$ is continuous, letting $\varepsilon\to 0$ yields $\Phi_{s}(u)=\Phi_{s}(v)$, as desired. ∎
Anhang A Proof of Proposition˜1
We establish the proposition by assembling a collection of lemmata.
Lemma 1.
Assume that $2\leq\mathopen{}\mathclose{{\left\lvert\mathcal{S}}}\right\rvert$$\infty$, and consider a family $(\phi_{s})_{s\in\mathcal{S}}$ of functions $\mathbf{R}\to\mathbf{R}$.
  1. (a)
    Property (1enumi) holds if either (i) for each $s\in\mathcal{S}$, $\phi_{s}$ is concave and weakly increasing or (ii) for some $s\in\mathcal{S}$, $\phi_{s}$ is constant.
  2. (b)
    Property (2enumi) holds if either (i) for each $s\in\mathcal{S}$, $\phi_{s}$ is concave and strictly increasing or (ii) for every $s\in\mathcal{S}$, $\phi_{s}$ is constant.
Proof of Lemma˜1.
Fix $X$ and $U$. For (a), note first that if $\phi_{s}$ is constant for some $s\in\mathcal{S}$, then
$\text{NSD}(X,\mathcal{S},U)\subseteq X=\text{NSD}(X,\mathcal{S},\Phi\circ U).$
Suppose instead that $\phi_{s}$ is concave and weakly increasing for each $s\in\mathcal{S}$. Fix any $x\notin\text{NSD}(X,\mathcal{S},\Phi\circ U)$; we will show that $x\notin\text{NSD}(X,\mathcal{S},U)$. By hypothesis, there is a mixture $p\in\Delta(X)$ such that
$\sum_{y\in X}\phi_{s}(U(y,s))p(y)>\phi_{s}(U(x,s)),$
for every $s\in\mathcal{S}$. Thus, for each $s\in\mathcal{S}$, since $\phi_{s}$ is concave
$\phi_{s}\mathopen{}\mathclose{{\left(\sum_{y\in X}U(y,s)p(y)}}\right)\geq\sum_{y\in X}\phi_{s}\mathopen{}\mathclose{{\left(U(y,s)}}\right)p(y)>\phi_{s}(U(x,s)),$
whence $\sum_{y\in X}U(y,s)p(y)>U(x,s)$ since $\phi_{s}$ is weakly increasing. [^10] A function $f:\mathbf{R}\to\mathbf{R}$ is weakly increasing if and only if $x\leq y$ implies $f(x)\leq f(y)$, which is equivalent (by contra-position) to ‘$f(x)>f(y)$ implies $x>y$’. It follows that $x\notin\text{NSD}(X,\mathcal{S},U)$.
For (b), note first that if $\phi_{s}$ is constant for every $s\in\mathcal{S}$, then
$\text{NWD}(X,\mathcal{S},U)\subseteq X=\text{NWD}(X,\mathcal{S},\Phi\circ U).$
Suppose instead that $\phi_{s}$ is concave and strictly increasing for every $s\in\mathcal{S}$. Fix any $x\notin\text{NWD}(X,\mathcal{S},\Phi\circ U)$; we will show that $x\notin\text{NWD}(X,\mathcal{S},U)$. By hypothesis, there is a mixture $p\in\Delta(X)$ such that
$\sum_{y\in X}\Phi(U(y,\cdot))p(y)>\Phi(U(x,\cdot)).$
For each $s\in\mathcal{S}$, since $\phi_{s}$ is concave and strictly increasing,
$\sum_{y\in X}\phi_{s}(U(y,s))p(y)\underset{(>)}{\geq}\phi_{s}(U(x,s))\quad\text{implies}\quad\sum_{y\in X}U(y,s)p(y)\underset{(>)}{\geq}U(x,s),$
by Pratt’s theorem. Hence $x\notin\text{NWD}(X,\mathcal{S},U)$. ∎
Our next lemma is purely technical and used in the final two lemmata.
Lemma 2.
For any $\varepsilon\in(0,1/2)$, any non-empty open interval $I\subseteq\mathbf{R}$, and any weakly increasing and non-constant $f:I\to\mathbf{R}$, there exist $u<v$ in $I$ and $\alpha\in[1/2,1/2+\varepsilon]$ such that $\alpha f(v)+(1-\alpha)f(u)>f(\frac{1}{2}u+\frac{1}{2}v)$.
Proof of Lemma˜2.
Suppose toward a contradiction that $f:I\to\mathbf{R}$ is weakly increasing and non-constant and that for some $\varepsilon\in(0,1/2)$, $\alpha f(v)+(1-\alpha)f(u)\leq f\mathopen{}\mathclose{{\left(\frac{1}{2}u+\frac{1}{2}v}}\right)$ for all $u<v$ in $I$ and $\alpha\in[1/2,1/2+\varepsilon]$. Since $f$ is Lebesgue-measurable (being weakly increasing), it follows by Sierpiński’s theorem (section˜2.2) that $f$ is concave. Note that $f$ is differentiable a.e. Choose any $m\in I$ at which $f$ is differentiable, with $f^{\prime}(m)>0$ (which must exist as $f$ is concave, weakly increasing, and non-constant), and observe that
$\lim_{\delta\to 0}\frac{\mathopen{}\mathclose{{\left(1/2+\varepsilon}}\right)f(m+\delta)+\mathopen{}\mathclose{{\left(1/2-\varepsilon}}\right)f(m-\delta)-f(m)}{\delta}=2\varepsilon f^{\prime}(m)>0.$
Hence, we may choose a (small) $\delta>0$ such that
$\mathopen{}\mathclose{{\left(\tfrac{1}{2}+\varepsilon}}\right)f(m+\delta)+\mathopen{}\mathclose{{\left(\tfrac{1}{2}-\varepsilon}}\right)f(m-\delta)>f(m)$
But then $\alpha f(v)+(1-\alpha)f(u)>f\mathopen{}\mathclose{{\left(\frac{1}{2}u+\frac{1}{2}v}}\right)$ holds for $\alpha\coloneqq 1/2+\varepsilon$, $u\coloneqq m-\delta$ and $v\coloneqq m+\delta$, a contradiction. ∎
Consider the following properties, which may or may not be satisfied jointly by a given collection $(I_{s})_{s\in\mathcal{S}}$ of non-empty open intervals of $\mathbf{R}$ and a transformation $\Phi:\mathbf{R}^{\mathcal{S}}\to\mathbf{R}^{\mathcal{S}}$:
  1. (1)
    $\text{NSD}(X,\mathcal{S},U)\subseteq\text{NSD}(X,\mathcal{S},\Phi\circ U)$ for every finite problem $(X,\mathcal{S},U)$ such that $U(X,s)\subseteq I_{s}$ for each $s\in\mathcal{S}$.
  2. (2)
    $\text{NWD}(X,\mathcal{S},U)\subseteq\text{NWD}(X,\mathcal{S},\Phi\circ U)$ for every finite problem $(X,\mathcal{S},U)$ such that $U(X,s)\subseteq I_{s}$ for each $s\in\mathcal{S}$.
(The arbitrary intervals $(I_{s})_{s\in\mathcal{S}}$ will not be required in this appendix, where only the special case $I_{s}=\mathbf{R}$ is relevant, but this extra generality will be utilised in the next appendix to prove Proposition˜2.)
Lemma 3.
Assume that $2\leq\mathopen{}\mathclose{{\left\lvert\mathcal{S}}}\right\rvert$$\infty$. For each $s\in\mathcal{S}$, fix a non-empty open interval $I_{s}\subseteq\mathbf{R}$ and a function $\phi_{s}:I_{s}\to\mathbf{R}$, and let $\Phi:\prod_{s\in\mathcal{S}}I_{s}\to\mathbf{R}^{\mathcal{S}}$ be $(\phi_{s}(u_{s}))_{s\in\mathcal{S}}$-situation-wise. Property (1) holds only if either (i) for each $s\in\mathcal{S}$, $\phi_{s}$ is concave and weakly increasing or (ii) for some $s\in\mathcal{S}$, $\phi_{s}$ is constant.
Beweis.
We assume that for all $s\in\mathcal{S}$, $\phi_{s}$ is not constant.
First, suppose that for some $i\in\mathcal{S}$, $\phi_{i}$ is not weakly increasing. Then, there exist $u_{i}>v_{i}$ in $I_{i}$ such that $\phi_{i}(u_{i})<\phi_{i}(v_{i})$. For each $j\in\mathcal{S}\setminus\{i\}$, since $\phi_{j}$ is not constant, there are $u_{j},v_{j}\in I_{j}$ such that $\phi_{j}(u_{j})<\phi_{j}(v_{j})$. This defines a pair of vectors in $\mathbf{R}^{\mathcal{S}}$, $u=(u_{j})_{j\in\mathcal{S}}$ and $v=(v_{j})_{j\in\mathcal{S}}$. Define $U:\{x,y\}\times\mathcal{S}\to\mathbf{R}$ by $U(x,\cdot)=u$ and $U(y,\cdot)=v$. Note that $U(\mathopen{}\mathclose{{\left\{x,y}}\right\},s)\subseteq I_{s}$ for each $s\in\mathcal{S}$, that $x\in\text{NSD}(\mathopen{}\mathclose{{\left\{x,y}}\right\},\mathcal{S},U)$ since $U(x,i)>U(y,i)$, and that $\text{NSD}(\mathopen{}\mathclose{{\left\{x,y}}\right\},\mathcal{S},\Phi\circ U)=\mathopen{}\mathclose{{\left\{y}}\right\}$. Hence $\text{NSD}(\mathopen{}\mathclose{{\left\{x,y}}\right\},\mathcal{S},U)\nsubseteq\text{NSD}(\mathopen{}\mathclose{{\left\{x,y}}\right\},\mathcal{S},\Phi\circ U)$.
Second, suppose that $\phi_{s}$ is weakly increasing for all $s\in\mathcal{S}$, and that $\phi_{i}$ is not concave for some $i\in\mathcal{S}$. Since $\phi_{i}$ is weakly increasing, it is Lebesgue-measurable, so by Sierpiński’s theorem (section˜2.2), there are $u_{i}<v_{i}$ in $I_{i}$ such that
$\tfrac{1}{2}\phi_{i}(u_{i})+\tfrac{1}{2}\phi_{i}(v_{i})>\phi_{i}\mathopen{}\mathclose{{\left(\tfrac{1}{2}u_{i}+\tfrac{1}{2}v_{i}}}\right).$
Fix an $\varepsilon>0$ small enough that
$\alpha\phi_{i}(u_{i})+(1-\alpha)\phi_{i}(v_{i})>\phi_{i}\mathopen{}\mathclose{{\left(\tfrac{1}{2}u_{i}+\tfrac{1}{2}v_{i}}}\right)$
for all $\alpha\in[1/2,1/2+\varepsilon]$. Fix another $j\in\mathcal{S}\setminus\{i\}$. Since $\phi_{j}$ is non-constant and weakly increasing, by Lemma˜2, there are $u_{j}<v_{j}$ in $I_{j}$ and an $\alpha\in[1/2,1/2+\varepsilon]$ such that
$\alpha\phi_{j}(v_{j})+(1-\alpha)\phi_{j}(u_{j})>\phi_{j}\mathopen{}\mathclose{{\left(\tfrac{1}{2}v_{j}+\tfrac{1}{2}u_{j}}}\right).$
For all $k\in\mathcal{S}\setminus\mathopen{}\mathclose{{\left\{i,j}}\right\}$, there exist $u_{k}<v_{k}$ in $I_{k}$ such that $\phi_{k}(u_{k})<\phi_{k}(v_{k})$ (as each $\phi_{k}$ is non-constant and weakly increasing). So, fix a three-element set $X\coloneqq\{x,y,z\}$, and consider the finite problem $(X,\mathcal{S},U)$ defined by, for all $k\in\mathcal{S}\setminus\mathopen{}\mathclose{{\left\{i,j}}\right\}$,
[…]
U(row,column) ijk
Note that $U(X,s)\subseteq I_{s}$ for each $s\in\mathcal{S}$, $y\in\text{NSD}(X,\mathcal{S},U)$, and the mixture $p\in\Delta(X)$ given by $p(x)\coloneqq\alpha$, $p(y)\coloneqq 0$ and $p(z)\coloneqq 1-\alpha$ satisfies $\sum_{w\in X}\Phi(U(w,\cdot))p(w)\gg\Phi(U(y,\cdot))$, so that $y\notin\text{NSD}(X,\mathcal{S},\Phi\circ U)$. ∎
Lemma 4.
Assume that $2\leq\mathopen{}\mathclose{{\left\lvert\mathcal{S}}}\right\rvert$$\infty$. For each $s\in\mathcal{S}$, fix a non-empty open interval $I_{s}\subseteq\mathbf{R}$ and a function $\phi_{s}:I_{s}\to\mathbf{R}$, and let $\Phi:\prod_{s\in\mathcal{S}}I_{s}\to\mathbf{R}^{\mathcal{S}}$ be $(\phi_{s})_{s\in\mathcal{S}}$-situation-wise. Property (2) holds only if either (i) for each $s\in\mathcal{S}$, $\phi_{s}$ is concave and strictly increasing or (ii) for every $s\in\mathcal{S}$, $\phi_{s}$ is constant.
Beweis.
We prove the contra-positive. Suppose there is at least one $i^{\star}\in\mathcal{S}$ such that $\phi_{i^{\star}}$ is not constant. We consider three cases:
  1. (1)
    Case 1: for some $i\in\mathcal{S}$, $\phi_{i}$ fails to be weakly increasing.
  2. (2)
    Case 2: case 1 does not hold and, for some $i\in\mathcal{S}$, $\phi_{i}$ fails to be strictly increasing.
  3. (3)
    Case 3: cases 1 and 2 do not hold and, for some $i\in\mathcal{S}$, $\phi_{i}$ fails to be concave.
Case 1: for some $i$, $\phi_{i}$ fails to be weakly increasing. Fix $u<v$ in $I_{i}$ such that $\phi_{i}(u)>\phi_{i}(v)$. Fix a two-element set $X\coloneqq\{x,y\}$, and consider a finite problem $(X,\mathcal{S},U)$ such that $U(X,s)\subseteq I_{s}$ for each $s\in\mathcal{S}$, $U(x,i)=u$, $U(y,i)=v$, and $U(x,j)=U(y,j)$ for each $j\in\mathcal{S}\setminus\{i\}$. Note that $\text{NWD}(X,\mathcal{S},U)=\{y\}\nsubseteq\{x\}=\text{NWD}(X,\mathcal{S},\Phi\circ U)$.
Case 2: not case 1, and for some $i$, $\phi_{i}$ fails to be strictly increasing. Recall that there is at least one $i^{\star}\in\mathcal{S}$ whose $\phi_{i^{\star}}$ is not constant.
  • If there exists $k\in\mathcal{S}\setminus\{i\}$ with non-constant $\phi_{k}$, then there exist $u_{i}<v_{i}$ in $I_{i}$ and $u_{k}<v_{k}$ in $I_{k}$ such that $\phi_{i}(u_{i})=\phi_{i}(v_{i})$ and $\phi_{k}(u_{k})<\phi_{k}(v_{k})$.
  • If instead $i$ is the only situation such that $\phi_{i}$ is not constant (i.e., $\phi_{j}$ is constant for every $j\in\mathcal{S}\setminus\{i\}$), then fixing an arbitrary other $j\in\mathcal{S}\setminus\{i\}$, there exist $u_{j}<v_{j}$ in $I_{j}$ and $u_{i}<v_{i}$ in $I_{i}$ such that $\phi_{j}(u_{j})=\phi_{j}(v_{j})$ and $\phi_{i}(u_{i})<\phi_{i}(v_{i})$.
Hence, either way, there exist situations $j\neq k$ in $\mathcal{S}$ and values $u_{j}<v_{j}$ in $I_{j}$ and $u_{k}<v_{k}$ in $I_{k}$ such that $\phi_{j}(u_{j})=\phi_{j}(v_{j})$ and $\phi_{k}(u_{k})<\phi_{k}(v_{k})$. Fix a two-element set $X\coloneqq\{x,y\}$, and consider a finite problem $(X,\mathcal{S},U)$ such that $U(X,s)\subseteq I_{s}$ for each $s\in\mathcal{S}$, $U(x,j)=u_{j}$, $U(y,j)=v_{j}$, $U(x,k)=v_{k}$, $U(y,k)=u_{k}$, and $U(x,\ell)=U(y,\ell)$ for each $\ell\in\mathcal{S}\setminus\{j,k\}$. Note that $\text{NWD}(X,\mathcal{S},U)=\{x,y\}\nsubseteq\{x\}=\text{NWD}(X,\mathcal{S},\Phi\circ U)$.
Case 3: not case 1 or case 2, and for some $i$, $\phi_{i}$ fails to be concave. Since $\phi_{i}$ is weakly increasing, it is Lebesgue-measurable, so by Sierpiński’s theorem (section˜2.2), there are $u_{i}<v_{i}$ in $I_{i}$ such that
$\tfrac{1}{2}\phi_{i}(u_{i})+\tfrac{1}{2}\phi_{i}(v_{i})>\phi_{i}\mathopen{}\mathclose{{\left(\tfrac{1}{2}u_{i}+\tfrac{1}{2}v_{i}}}\right).$
Fix an $\varepsilon>0$ small enough that
$\alpha\phi_{i}(u_{i})+(1-\alpha)\phi_{i}(v_{i})>\phi_{i}\mathopen{}\mathclose{{\left(\tfrac{1}{2}u_{i}+\tfrac{1}{2}v_{i}}}\right)$
for all $\alpha\in[1/2,1/2+\varepsilon]$. Fix another situation $j\in\mathcal{S}\setminus\{i\}$. Since $\phi_{j}$ is strictly increasing, by Lemma˜2, there are $u_{j}<v_{j}$ in $I_{j}$ and an $\alpha\in[1/2,1/2+\varepsilon]$ such that
$\alpha\phi_{j}(v_{j})+(1-\alpha)\phi_{j}(u_{j})>\phi_{j}\mathopen{}\mathclose{{\left(\tfrac{1}{2}v_{j}+\tfrac{1}{2}u_{j}}}\right).$
Fix a three-element set $X\coloneqq\{x,y,z\}$, and consider a finite problem $(X,\mathcal{S},U)$ such that $U(X,s)\subseteq I_{s}$ for each $s\in\mathcal{S}$, $U(\cdot,k)$ is constant for all $k\in\mathcal{S}\setminus\{i,j\}$, and
[…]
U(row,column) ij
Note that $y\in\text{NWD}(X,\mathcal{S},U)$, and the mixture $p\in\Delta(X)$ given by $p(x)\coloneqq\alpha$, $p(y)\coloneqq 0$ and $p(z)\coloneqq 1-\alpha$ satisfies $\sum_{w\in X}\Phi(U(w,\cdot))p(w)>\Phi(U(y,\cdot))$, so that $y\notin\text{NWD}(X,\mathcal{S},\Phi\circ U)$. ∎
Proof of Proposition˜1.
By Lemma˜1, the conditions in (b) and (c) are sufficient for, respectively, property (1enumi) and property (2enumi). By Lemma˜3, the condition in (b) is necessary for property (1enumi), and by Lemma˜4, the condition in (c) is necessary for property (2enumi).
It remains to derive the characterisation of property (1) by that of property (1enumi) and to show that property (2enumi) is equivalent to property (2). To that end, note that by the above, property (1) holds if and only if
  1. (i)
    $\phi_{s}$ is weakly increasing for all $s\in\mathcal{S}$ and, for any $S\subseteq\mathcal{S}$ with $\mathopen{}\mathclose{{\left\lvert S}}\right\rvert\geq 2$, $(\phi_{s})_{s\in S}$ satisfies the conditions stated in Proposition˜1(b)
(where the ‘only if’ direction follows by applying property (1) to finite problems, $(X,S,U)$, for each non-empty subset $S\subseteq\mathcal{S}$ of situations), that property (2) holds if
  1. (ii)
    $\phi_{s}$ is weakly increasing for all $s\in\mathcal{S}$ and, for any $S\subseteq\mathcal{S}$ with $\mathopen{}\mathclose{{\left\lvert S}}\right\rvert\geq 2$, $(\phi_{s})_{s\in S}$ satisfies the conditions stated in Proposition˜1(c),
and that property (2) holds only if
  1. (iii)
    for any $S\subseteq\mathcal{S}$ with $\mathopen{}\mathclose{{\left\lvert S}}\right\rvert\geq 2$, $(\phi_{s})_{s\in S}$ satisfies the conditions stated in Proposition˜1(c).
This concludes the proof since (i) is equivalent to $(\phi_{s})_{s\in\mathcal{S}}$ satisfying the properties stated in Proposition˜1(a), and both (ii) and (iii) are equivalent to $(\phi_{s})_{s\in\mathcal{S}}$ satisfying the properties stated in Proposition˜1(c). ∎
Anhang B Proof of Proposition˜2
We establish the proposition by assembling a collection of lemmata.
Lemma 5.
Assume that $2\leq\mathopen{}\mathclose{{\left\lvert\mathcal{S}}}\right\rvert$$\infty$. Fix a family $(\phi_{s})_{s\in\mathcal{S}}$ of functions $\mathbf{R}\to\mathbf{R}$, and let $\Phi:\mathbf{R}^{\mathcal{S}}\to\mathbf{R}^{\mathcal{S}}$ be $(\phi_{s})_{s\in\mathcal{S}}$-situation-wise. Properties (3) and (4) hold if for each $s\in\mathcal{S}$, $\phi_{s}$ is convex and strictly increasing.
Beweis.
Suppose that for every $s\in\mathcal{S}$, $\phi_{s}$ is convex and strictly increasing. Fix any $x\notin\text{NWD}(X,S,U)$ ($x\notin\text{NSD}(X,S,U)$); we will show that $x\notin\text{NWD}(X,S,\Phi\circ U)$ ($x\notin\text{NSD}(X,S,\Phi\circ U)$). By hypothesis, there is a mixture $p\in\Delta(X)$ such that
$\sum_{y\in X}U(y,\cdot)p(y)\underset{(\gg)}{>}U(x,\cdot).$
By Pratt’s theorem, it must, therefore, be that
$\sum_{y\in X}\Phi(U(y,\cdot))p(y)\underset{(\gg)}{>}\Phi(U(x,\cdot)),$
as desired.∎
Lemma 6.
Assume that $2\leq\mathopen{}\mathclose{{\left\lvert\mathcal{S}}}\right\rvert$$\infty$. Fix a family $(\phi_{s})_{s\in\mathcal{S}}$ of functions $\mathbf{R}\to\mathbf{R}$, and let $\Phi:\mathbf{R}^{\mathcal{S}}\to\mathbf{R}^{\mathcal{S}}$ be $(\phi_{s})_{s\in\mathcal{S}}$-situation-wise. If either property (3enumi) or property (4enumi) holds, then for each $s\in\mathcal{S}$, $\phi_{s}:\mathbf{R}\to\mathbf{R}$ is strictly increasing and continuous.
Beweis.
Suppose that either property (3enumi) or property (4enumi) holds. Then, it is immediate that each $\phi_{s}$ must be strictly increasing: if for some $s\in\mathcal{S}$ and some $a<b$, $\phi_{s}(a)\geq\phi_{s}(b)$, it is easy to construct a problem where some $x$ is dominated (either weakly or strictly) pre-transformation and undominated after, due to a strict inequality in $s$ failing after the transformation.
For all $u\in\mathbf{R}$, let
$\phi_{s}(u-)=\sup_{(-\infty,u)}\phi_{s}\quad\text{and}\quad\phi_{s}(u+)=\inf_{(u,\infty)}\phi_{s}.$
Suppose that $\phi_{s}$ is discontinuous at $u\in\mathbf{R}$ for some $s\in\mathcal{S}$ and seek a contradiction. It suffices to exhibit $a,b,c\in\mathbf{R}^{\mathcal{S}}$ such that $c\notin\mathrm{NSD}(\{a,b,c\},\mathcal{S},\mathrm{id})$ and $c\in\mathrm{NWD}(\{a,b,c\},\mathcal{S},\Phi)$, where $\mathrm{id}(x,\cdot)\coloneqq x$ for each $x\in\{a,b,c\}$. To this end, note that
$\phi_{s}(u+)>\frac{\phi_{s}(u-)+\phi_{s}(v)}{2}$ (1)
for $v>u$ sufficiently close to $u$, since $\phi_{s}$ is strictly increasing and discontinuous at $u$. Fix $t\in\mathcal{S}\setminus\{s\}$ and note that $\phi_{t}$ has countably many discontinuities since it is strictly increasing. Then we may choose $v>u$ such that (1) holds and $\phi_{t}$ is continuous at $v$. Since $\phi_{t}$ is continuous at $v$ and strictly increasing, there exists $w\in(u,v)$ such that
$\phi_{t}(w-)>\frac{\phi_{t}(u+)+\phi_{t}(v)}{2}.$ (2)
Given $v^{\prime}<u$ and $w^{\prime}\in(u,w)$, let $a,b,c\in\mathbf{R}^{\mathcal{S}}$ be given by
$a_{r}=\begin{cases}v&\text{if $r=s$}\\ v^{\prime}&\text{if $r\neq s$}\end{cases}\qquad b_{r}=\begin{cases}v^{\prime}&\text{if $r=s$}\\ v&\text{if $r\neq s$.}\end{cases}\qquad c_{r}=\begin{cases}w^{\prime}&\text{if $r=s$}\\ w&\text{if $r\neq s$.}\end{cases}$
Since $\max\{u,w\}<v$, we may choose $v^{\prime}$ and $w^{\prime}$ sufficiently close to $u$ to ensure that $c\notin\mathrm{NSD}(\{a,b,c\},\mathcal{S},\mathrm{id})$.
It remains to show that $c\in\mathrm{NWD}(\{a,b,c\},\mathcal{S},\Phi)$. To this end, note that $b_{s}<c_{s}<a_{s}$ and $a_{t}<c_{t}<b_{t}$, so that $\Phi_{s}(b)<\Phi_{s}(c)<\Phi_{s}(a)$ and $\Phi_{t}(a)<\Phi_{t}(c)<\Phi_{t}(b)$ (since $\Phi$ is situation-wise and each $\phi_{r}$ is strictly increasing). Then, it suffices to show that
$\Phi_{r}(c)>\frac{\Phi_{r}(a)+\Phi_{r}(b)}{2}$
for each $r\in\{s,t\}$. For the case $r=s$, note that
$\Phi_{s}(c)=\phi_{s}(w^{\prime})\geq\phi_{s}(u+)>\frac{\phi_{s}(u-)+\phi_{s}(v)}{2}$ $\geq\frac{\phi_{s}(v^{\prime})+\phi_{s}(v)}{2}$
$=\frac{\Phi_{s}(b)+\Phi_{s}(a)}{2},$
where the first inequality holds since $w^{\prime}>u$, the second is (1), and the third holds since $u>v^{\prime}$. For the case $r=t$, note that
$\Phi_{t}(c)=\phi_{t}(w)\geq\phi_{t}(w-)>\frac{\phi_{t}(u+)+\phi_{t}(v)}{2}$ $\geq\frac{\phi_{t}(v^{\prime})+\phi_{t}(v)}{2}$
$=\frac{\Phi_{t}(a)+\Phi_{t}(b)}{2},$
where the first inequality holds by monotonicity, the second is (2), and the third holds since $u>v^{\prime}$.
Accordingly, $c\in\mathrm{NWD}(\{a,b,c\},\mathcal{S},\Phi)\subseteq\mathrm{NSD}(\{a,b,c\},\mathcal{S},\Phi)$ but $c\notin\mathrm{NSD}(\{a,b,c\},\mathcal{S},\mathrm{id})\supseteq\mathrm{NWD}(\{a,b,c\},\mathcal{S},\mathrm{id})$, contradicting properties (3enumi) and (4enumi). ∎
Lemma 7.
Assume that $2\leq\mathopen{}\mathclose{{\left\lvert\mathcal{S}}}\right\rvert$$\infty$. Fix a family $(\phi_{s})_{s\in\mathcal{S}}$ of functions $\mathbf{R}\to\mathbf{R}$, and let $\Phi:\mathbf{R}^{\mathcal{S}}\to\mathbf{R}^{\mathcal{S}}$ be $(\phi_{s})_{s\in\mathcal{S}}$-situation-wise. If either property (3enumi) or property (4enumi) holds, then for each $s\in\mathcal{S}$, $\phi_{s}:\mathbf{R}\to\mathbf{R}$ is strictly increasing and convex.
Beweis.
Suppose that either property (3enumi) or property (4enumi) holds. Then, by Lemma 6, each $\phi_{s}$ must be strictly increasing and continuous.
To establish convexity, suppose first that property (3enumi) holds. Since each $\phi_{s}$ is strictly increasing and continuous, each image $I_{s}\coloneqq\phi_{s}(\mathbf{R})$ is a (non-empty) open interval of $\mathbf{R}$, and $\phi_{s}$ admits an inverse $\phi_{s}^{-1}:I_{s}\to\mathbf{R}$. Since $\Phi$ is $(\phi_{s})_{s\in\mathcal{S}}$-situation-wise, it admits an inverse $\Phi^{-1}:I\to\mathbf{R}^{\mathcal{S}}$, where $I\coloneqq\prod_{s\in\mathcal{S}}I_{s}$. Note that, given any finite problem $(X,\mathcal{S},V)$ with $V(X)\subseteq I$, $\mathrm{NSD}(X,\mathcal{S},V)\subseteq\mathrm{NSD}(X,\mathcal{S},\Phi^{-1}\circ V)$. Then, Lemma˜3 implies that each $\phi_{s}^{-1}$ is concave; thus, each $\phi_{s}$ is convex.
The same steps, with Lemma˜4 used in place of Lemma˜3, deliver convexity of each $\phi_{s}$ for the case in which property (4enumi) holds. ∎
Proof of Proposition˜2.
By Lemma˜5, if $\phi_{s}$ is convex and strictly increasing for each $s\in\mathcal{S}$, then properties (3) and (4) hold, which implies that properties (3enumi) and (4enumi) hold. The equivalence is closed by Lemma 7. ∎
Anhang C Proof of Theorem˜1
In light of Propositions˜1 and 2, what must be shown is that $\Phi$ being situation-wise is necessary for expansion and for contraction of NSD and of NWD. We do this via two lemmata.
Lemma 8.
Assume that $2\leq\mathopen{}\mathclose{{\left\lvert\mathcal{S}}}\right\rvert$$\infty$, and consider a transformation $\Phi:\mathbf{R}^{\mathcal{S}}\to\mathbf{R}^{\mathcal{S}}$. If property (1) or property (2) holds, then $\Phi$ is situation-wise.
Proof of Lemma˜8.
Fix $s\in\mathcal{S}$ and take any $u,v\in\mathbf{R}^{\mathcal{S}}$ with $u_{s}=v_{s}$. Consider the finite problem $(\{x,y\},\{s\},U)$, where $x\neq y$, $U(x,\cdot)\coloneqq u$ and $U(y,\cdot)\coloneqq v$. Since $u_{s}=v_{s}$, neither object weakly dominates the other, so $\{x,y\}=\text{NWD}(\{x,y\},\{s\},U)=\text{NSD}(\{x,y\},\{s\},U)$. Assuming that property (2) (property (1)) holds, $\{x,y\}=\text{NWD}(\{x,y\},\{s\},\Phi\circ U)$ ($\{x,y\}=\text{NSD}(\{x,y\},\{s\},\Phi\circ U)$) as well. Then $\Phi_{s}(u)=\Phi_{s}(v)$, or else the object with the larger $s$-coordinate after the transformation strictly dominates the other on $S=\{s\}$. Since $\Phi_{s}(u)=\Phi_{s}(v)$ for all $u,v\in\mathbf{R}^{\mathcal{S}}$ with $u_{s}=v_{s}$, there exists a function $\phi_{s}:\mathbf{R}\to\mathbf{R}$ with $\Phi_{s}(w)=\phi_{s}(w_{s})$ for all $w$. Since $s$ was arbitrary, $\Phi$ is situation-wise. ∎
Lemma 9.
Assume that $2\leq\mathopen{}\mathclose{{\left\lvert\mathcal{S}}}\right\rvert$$\infty$, and consider a transformation $\Phi:\mathbf{R}^{\mathcal{S}}\to\mathbf{R}^{\mathcal{S}}$. If property (3) or property (4) holds, then there exists a family $(\phi_{s})_{s\in\mathcal{S}}$ of continuous and strictly increasing functions $\mathbf{R}\to\mathbf{R}$ such that $\Phi$ is $(\phi_{s})_{s\in\mathcal{S}}$-situation-wise.
Proof of Lemma˜9.
For each $s\in\mathcal{S}$, let $\pi^{s}:\mathbf{R}\to\mathbf{R}^{S}$ be given by $\pi^{s}_{s}(u)=u$ and $\pi^{s}_{t}(u)=0$ for all $t\neq s$, and set $\phi_{s}=\Phi_{s}\circ\pi^{s}$. Recall that for all $u\in\mathbf{R}$,
$\phi_{s}(u-)=\sup_{(-\infty,u)}\phi_{s}\quad\text{and}\quad\phi_{s}(u+)=\inf_{(u,\infty)}\phi_{s}.$
Note that, for all $a,b\in\mathbf{R}^{\mathcal{S}}$ such that $a_{s}>b_{s}$, $\Phi_{s}(a)>\Phi_{s}(b)$. This implies that $\phi_{s}$ is strictly increasing and that $\phi_{s}(a_{s}-\varepsilon)<\Phi_{s}(a)<\phi_{s}(a_{s}+\varepsilon)$ for all $a\in\mathbf{R}^{\mathcal{S}}$ and $\varepsilon>0$. Letting $\varepsilon$ vanish yields
$\phi_{s}(a_{s}-)\leq\Phi_{s}(a)\leq\phi_{s}(a_{s}+)\quad\text{for all $a\in\mathbf{R}^{\mathcal{S}}$.}$ (3)
Then, it is enough to show that $\phi_{s}$ is continuous for each $s\in\mathcal{S}$, as this implies that $\Phi_{s}(a)=\phi_{s}(a_{s})$ for all $a\in\mathbf{R}^{\mathcal{S}}$.
We now mimic the proof of Lemma˜6 nearly verbatim. Suppose that $\phi_{s}$ is discontinuous at $u\in\mathbf{R}$ for some $s\in\mathcal{S}$ and seek a contradiction. It suffices to exhibit $a,b,c\in\mathbf{R}^{\mathcal{S}}$ such that $c\notin\mathrm{NSD}(\{a,b,c\},\mathcal{S},\mathrm{id})$ and $c\in\mathrm{NWD}(\{a,b,c\},\mathcal{S},\Phi)$, where $\mathrm{id}(x,\cdot)\coloneqq x$ for each $x\in\{a,b,c\}$. To this end, note that
$\phi_{s}(u+)>\frac{\phi_{s}(u-)+\phi_{s}(v)}{2}$ (4)
for $v>u$ sufficiently close to $u$, since $\phi_{s}$ is strictly increasing and discontinuous at $u$. Fix $t\in\mathcal{S}\setminus\{s\}$ and note that $\phi_{t}$ has countably many discontinuities since it is strictly increasing. Then we may choose $v>u$ such that (4) holds and $\phi_{t}$ is continuous at $v$. Since $\phi_{t}$ is continuous at $v$ and strictly increasing, there exists $w\in(u,v)$ such that
$\phi_{t}(w-)>\frac{\phi_{t}(u+)+\phi_{t}(v)}{2}.$ (5)
Given $v^{\prime}<u$ and $w^{\prime}\in(u,w)$, let $a,b,c\in\mathbf{R}^{\mathcal{S}}$ be given by
$a_{r}=\begin{cases}v&\text{if $r=s$}\\ v^{\prime}&\text{if $r\neq s$}\end{cases}\qquad b_{r}=\begin{cases}v^{\prime}&\text{if $r=s$}\\ v&\text{if $r\neq s$.}\end{cases}\qquad c_{r}=\begin{cases}w^{\prime}&\text{if $r=s$}\\ w&\text{if $r\neq s$.}\end{cases}$
Since $\max\{u,w\}<v$, we may choose $v^{\prime}$ and $w^{\prime}$ sufficiently close to $u$ to ensure that $c\notin\mathrm{NSD}(\{a,b,c\},\mathcal{S},\mathrm{id})$.
It remains to show that $c\in\mathrm{NWD}(\{a,b,c\},\mathcal{S},\Phi)$. To this end, note that $b_{s}<c_{s}<a_{s}$ and $a_{t}<c_{t}<b_{t}$, so that $\Phi_{s}(b)<\Phi_{s}(c)<\Phi_{s}(a)$ and $\Phi_{t}(a)<\Phi_{t}(c)<\Phi_{t}(b)$, from above. Then, it suffices to show that
$\Phi_{r}(c)>\frac{\Phi_{r}(a)+\Phi_{r}(b)}{2}$
for each $r\in\{s,t\}$. For the case $r=s$, note that
$\Phi_{s}(c)\geq\phi_{s}(w^{\prime}-)\geq\phi_{s}(u+)>\frac{\phi_{s}(u-)+\phi_{s}(v)}{2}$ $\geq\frac{\phi_{s}(v^{\prime})+\phi_{s}(v)}{2}$
$=\frac{\Phi_{s}(b)+\Phi_{s}(a)}{2},$
where the first inequality and the equality follow from (3), the second inequality holds since $w^{\prime}>u$, the third is (4), and the fourth holds since $u>v^{\prime}$. For the case $r=t$, note that
$\Phi_{t}(c)\geq\phi_{t}(w-)>\frac{\phi_{t}(u+)+\phi_{t}(v)}{2}\geq\frac{\phi_{t}(v^{\prime})+\phi_{t}(v)}{2}=\frac{\Phi_{t}(a)+\Phi_{t}(b)}{2}$
where the first inequality and the equality follow from (3), the second inequality is (5), and the third holds since $u>v^{\prime}$. ∎
Proof of Theorem˜1.
Propositions˜1 and 2 establish sufficiency direction. For necessity, in light of Propositions˜1 and 2, it suffices to show that $\Phi$ being situation-wise is necessary for each of properties (1)(4), and that follows from Lemmata˜9 and 8. ∎
Anhang D Proof of Theorem˜2
Proof of Theorem˜2.
The ‘if’ parts of Theorem˜2(a)(b) are easily established.
For the ‘only if’ parts, mimicking Theorem˜1’s proof exactly, we conclude that $\Phi$ being $(\phi_{s})_{s\in\mathcal{S}}$-situation-wise (for some family $(\phi_{s})_{s\in\mathcal{S}}$ of maps $\mathbf{R}\to\mathbf{R}$) is necessary for both properties ($\hskip 1.00006pt\widehat{1}\hskip 1.00006pt$) and ($\hskip 1.00006pt\widehat{2}\hskip 1.00006pt$) to hold. We now argue that $\phi_{s}$ being weakly increasing for each $s\in\mathcal{S}$ is necessary as well. Suppose that $\Phi$ is $(\phi_{s})_{s\in\mathcal{S}}$-situation-wise for some family $(\phi_{s})_{s\in\mathcal{S}}$, but there are $s\in\mathcal{S}$ and $u_{s}<v_{s}$ in $\mathbf{R}$ such that $\phi_{s}(u_{s})>\phi_{s}(v_{s})$, and consider a finite problem $(\{x,y\},\mathcal{S},U)$ with $U$ such that $U(x,s)=u_{s}$ and $U(y,s)=v_{s}$. Note that
$\widehat{\text{NSD}}(X,\{s\},U)=\{y\}=\widehat{\text{NWD}}(X,\{s\},U)$
and that
$\widehat{\text{NSD}}(X,\{s\},\Phi\circ U)=\{x\}=\widehat{\text{NWD}}(X,\{s\},\Phi\circ U),$
so that properties ($\hskip 1.00006pt\widehat{1}\hskip 1.00006pt$) and ($\hskip 1.00006pt\widehat{2}\hskip 1.00006pt$) both fail.
[…]
The ‘if’ parts of Theorem˜2(c)(d) are immediate. For the ‘only if’ parts, note first that, if $u_{s}>v_{s}$ and $\Phi_{s}(u)\leq\Phi_{s}(v)$ for some $u,v\in\mathbf{R}^{\mathcal{S}}$ and $s\in\mathcal{S}$, then Properties ($\hskip 1.00006pt\widehat{3}\hskip 1.00006pt$) and ($\hskip 1.00006pt\widehat{4}\hskip 1.00006pt$) fail for the finite problem $(\{x,y\},\{s\},U)$ where $U(x,\cdot)\coloneqq u$ and $U(y,\cdot)\coloneqq v$. The ‘only if’ part of Theorem˜2(c) follows, and to establish the ‘only if’ part of Theorem˜2(d), it suffices to show that property ($\hskip 1.00006pt\widehat{4}\hskip 1.00006pt$) fails if, for some $u,v\in\mathbf{R}^{\mathcal{S}}$ and $s,t\in\mathcal{S}$, $u_{s}=v_{s}$, $\Phi_{s}(u)>\Phi_{s}(v)$, and $u_{t}<v_{t}$; to see this, consider the finite problem $(\{x,y\},\{s,t\},U)$ where $U(x,\cdot)\coloneqq u$ and $U(y,\cdot)\coloneqq v$. ∎